OpenStudy (anonymous):

If (x + 2) is a factor of x^3 − 6x^2 + kx + 10, k =?

OpenStudy (mathmath333):

\(\large \begin{align} \color{black}{ \normalsize \text{if (x+2) if a factor of f(x) then,} \hspace{.33em}\\~\\ f(-2)=0 \hspace{.33em}\\~\\ }\end{align}\)

OpenStudy (xapproachesinfinity):

yep :)

OpenStudy (anonymous):

MInd blown kaboom.

OpenStudy (mathmath333):

\(\large \begin{align} \color{black}{ f(x)=x^3-6x^2+kx+10 \hspace{.33em}\\~\\ f(-2)=(-2)^3-6(-2)^2+k(-2)+10=0 \hspace{.33em}\\~\\ }\end{align}\)

OpenStudy (anonymous):

Smh. I should have know that. One sec.

OpenStudy (anonymous):